Agenda/Minutes on 2020/9/25
CRY+VIS Interface Meeting
Participants: Takahashi, Miyo, Uchiyama, Aso, Hirata, Sato, Ikeda, Tomaru, Hagiwara, Kimura, Ushiba, Sawada、Akutsu
- The schedule reconstruction for vacuum chamber opening is necessary due to the extension of RSE Trial.
- Removing of SR3 is planed from end of January with moving of the assembly frame.
- Confirmation of timing for the tower-payload disconnection.
- Request on the IP tracking at evacuation and floor monitor at vent in IYV.
Participants: Takahashi, Fabian, Hirata, Satou, Terrence, Miyo, Aso, Washimi, Ishizaki
- Iguchi-san will join VIS.
- [Fujii] Can you please update the status of paper writing especially about the Type-A mechanics and its control? I also would like to know the rough schedule that you team are planning. Please let me know. Leaving your comments here will be enough. O-tsu-ka-re-sa-ma-de-su. (My account will have to be removed I guess btw..)
[Sato] Design of the limit switches for BS IP FR. BS IP FR limit switch_20200924.pdf
I submitted the paper to C&QG, finally.
- Hirata-san and I had an information session for Washimi-san.
- I asked to Takahashi-san for some long poles to mount the dial gauges around outer frames. Maybe there are some in TAMA. Work in progress.
- Given the new temperature conditions in the central room, I offloaded fishing rods:
BS: F0 and BF fishing rods look healthy and F1 one is at the bottom of its range (klog 15071).
SR3: F0 fishig rod is healthy, the F1 one is still jammed and BF one at its upper limit (klog 15070).
- I changed some DC filters setpoint and updated BURT and SDF files. In SR3 I keep the height of the mirror constant.
For SR3 F1, I calculated the amount of ballast the needs to be removed to bring it to nominal height: 31 grams or less (klog 15102). This number has be reviewed once I calculate how much mass we need to remove from the IM. The BF fishing rod needs to be in mid-range with control off to know.
- Next week:
- Same calculation for the other filters: they require setting the fishing rods to mid-range.
- With the new temperature conditions I need to measure IP resonant frequencies again.
- LVDT card document
- Ordered the manufacture of a modified IMMT2 support frame.
- KAGRA is in "RSE Trial" now. "SEO is thinking about possibility for one week extension and the schedule reconstruction for vacuum chamber opening."
- Preparation work.
- Procedure documents.
Criteria in the acceptance test Sheet.
- PAB (10/8)
- Chief Meeting (10/12)
[Terrence] VIS health checking system VISHack
The week from 9/28
- Takahashi: 9/30-10/2
- Sato: 9/30-10/2
- Hirata: 9/30-10/2
The week from 10/5
- Takahashi: 10/5-9
- Sato: 10/5-9
- Hirata: 10/5-9
- On 2020/10/2(Fri)